OpenStudy (anonymous):

Let f(x) = 9x − 2 and g(x) = −x + 3. Find f(g(x)).

OpenStudy (anonymous):

@dawsonj

OpenStudy (dawsonj):

Let f(x) = 9x – 2 and g(x) = –x + 3. Find ... Replace the value of x in g(x) with 9x - 2 What are you getting? ... 1. f(g(x))=9(-x+3)-2 f(g(x))=-9x+27-2 f(g(x))=-9x+25.

OpenStudy (anonymous):

A)-9x-2 B)-9x+5 C)-9x+25 D)-9x+27

OpenStudy (anonymous):

Thank you @dawsonj

OpenStudy (dawsonj):

c

OpenStudy (anonymous):

f(g(x))=9(-x+3)-2 f(g(x))=-9x+27-2 How'd you get this part ?

OpenStudy (dawsonj):

am i right

OpenStudy (anonymous):

I'm pretty sure. It looks right to me, I just don't understand how you put the equation together which is where my problem lies.

OpenStudy (dawsonj):

i have the same problem

OpenStudy (anonymous):

oh xD So, how'd you figure it out ?

OpenStudy (dawsonj):

i did it i my head

OpenStudy (anonymous):

ahh, well as long as i have the answer i don't care lol. Thanks for your help (:

OpenStudy (dawsonj):

np

OpenStudy (damonte212):

..